1 The Bahasa language

The file bahasa.dtx^{1} defines all the language definition macros for the bahasa indonesia / bahasa melayu language. Bahasa just means ‘language’ in bahasa indonesia / bahasa melayu. Since both national versions of the language use the same writing, although differing in pronounciation, this file can be used for both languages.

For this language currently no special definitions are needed or available.

The macro \LdfInit takes care of preventing that this file is loaded more than once, checking the category code of the @ sign, etc.

1 ⟨∗code⟩

 2 \LdfInit{bahasa}\captionsbahasa

When this file is read as an option, i.e. by the \usepackage command, bahasa could be an ‘unknown’ language in which case we have to make it known. So we check for the existence of \l@bahasa to see whether we have to do something here.

 3 \ifx\l@bahasa\@undefined
 4 \@nopatterns{Bahasa}
 5 \adddialect\l@bahasa0\fi

The next step consists of defining commands to switch to (and from) the Bahasa language.

\captionsbahasa The macro \captionsbahasa defines all strings used in the four standard documentclasses provided with L^{A}TEX.

 6 \addto\captionsbahasa{%
 7 \def\prefacename{Pendahuluan}%
 8 \def\refname{Pustaka}%
 9 \def\abstractname{Ringkasan}% (sometime it's called 'intisari'
10  %  or 'ikhtisar')
11 \def\bibname{Bibliografi}%
12 \def\chaptername{Bab}%
13 \def\appendixname{Lampiran}%
14 \def\contentsname{Daftar Isi}%
15 \def\listfigurename{Daftar Gambar}%
16 \def\listtablename{Daftar Tabel}%
17 % Glossary: Daftar Istilah
18 \def\indexname{Indeks}%
19 \def\figurename{Gambar}%
20 \def\tablename{Tabel}%
21 \def\partname{Bagian}%
22 %  Subject:  Subyek
23 %  From:  Dari
24 \def\enclname{Lampiran}%
25 \def\ccname{cc}%
26 \def\headtoname{Kepada}%
27 \def\pagename{Halaman}%
28 %  Notes (Endnotes): Catatan
29 \def\seename{lihat}%
30 \def\alsoname{lihat juga}%
31 \def\proofname{Bukti}%

1The file described in this section has version number v1.0e and was last revised on 1996/12/23.

\datebahasa The macro \datebahasa redefines the command \today to produce Bahasa dates.

33 \def\datebahasa{%
34 \def\today{\number\day~\ifcase\month\or
35  Januari\or Februari\or Maret\or April\or Mei\or Juni\or
36  Juli\or Agustus\or September\or Oktober\or Nopember\or Desember\fi
37  \space \number\year}}

\extrasbahasa \noextrasbahasa The macro \extrasbahasa will perform all the extra definitions needed for the Bahasa language. The macro \extrasbahasa is used to cancel the actions of \extrasbahasa. For the moment these macros are empty but they are defined for compatibility with the other language definition files.

38 \addto\extrasbahasa{}
39 \addto\noextrasbahasa{}

\bahasahyphenmins The bahasa hyphenation patterns should be used with \lefthyphenmin set to 2

and \righthyphenmin set to 2.
40 \def\bahasahyphenmins{\tw@\tw@}

The macro \ldf@finish takes care of looking for a configuration file, setting the main language to be switched on at \begin{document} and resetting the category code of @ to its original value.

41 \ldf@finish{bahasa}

42 ⟨/code⟩